OpenStudy (anonymous):

(x-6)(x-2)-x(x-3)=(x-6)

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

omg... steps please? -___-

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

idk i used myalgebra.com

OpenStudy (anonymous):

wow.. cheap. this is why I like to LEARN the steps instead of just finding the answer ...

OpenStudy (anonymous):

well my bad

OpenStudy (anonymous):

you can do it in 2 ways 1.way expand and reduce the left handside,then move everything to one side making the other 0,factorize and solve

OpenStudy (asnaseer):

notice there is a term involving (x-6) on both sides of this equation.

OpenStudy (asnaseer):

if you move the (x-6) from the right-hand-side to the left-hand-side, then you can factorise it with the other term involving (x-6)

OpenStudy (asnaseer):

try doing this and see what you get

jimthompson5910 (jim_thompson5910):

(x-6)(x-2)-x(x-3)=(x-6) x^2 - 8x + 12-x(x-3)=(x-6) x^2 - 8x + 12-x^2+3x=x-6 -5x + 12 = x-6 Can you take it from here?